40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 17 23:28:32 2011, rana, Update, Electronics, StochMon 
    Reply  Wed Oct 19 18:18:03 2011, Suresh, Update, Electronics, StochMon : Filters installed PA190146.JPGPA190143.JPG
Message ID: 5682     Entry time: Mon Oct 17 23:28:32 2011     Reply to this: 5706
Author: rana 
Type: Update 
Category: Electronics 
Subject: StochMon 

To get to the bottom of the RFAM mystery, we've got to resurrect the StochMon to trend the RFAM after the IMC.

We will put an 1811 on the MC_TRANS or IP_POS beam (the 1811 has an input noise of 2.5 pW/rHz).

Then the Stochmon has an input pre-amp, some crappy filters, and then Wenzel RMS->DC converters. We will replace the hand-made filters with the following ones from Mini-Circuits which happen to match our modulation frequencies perfectly:

11 MHz     SBP-10.7+

55 MHz     SBP-60+

29.5 MHz   SBP-30+

ELOG V3.1.3-